OpenStudy (anonymous):

Express 6x^3+23x^2-6x-8 in factored form

hartnn (hartnn):

for the 1st root we will need guess and check...or is anything else, like hint is given ?

hartnn (hartnn):

after we find 1st root, its just the matter of synthetic division

OpenStudy (anonymous):

Um okay

hartnn (hartnn):

like check for x=1,-1,2,-2,3,-3,4,-4.... if for any of these, f(x) comes out to be 0, then we get our 1st root.

OpenStudy (anonymous):

Isn't there an easier way to do it, like a faster way ?

hartnn (hartnn):

yes, there is a theorem, rational root theorem, heard of it ?

hartnn (hartnn):

it simplifies guessing of a root

OpenStudy (anonymous):

Oh could you show me how to do it then?

hartnn (hartnn):

sure, if \( f(x)=ax^3+bx^2+cx+d \) then a possible rational root will be reduced form of \(\huge \pm \dfrac{d}{a}\)

hartnn (hartnn):

plus or minus , last term (constant) divided by co-efficient of highest power of x

hartnn (hartnn):

so, what can you guess about the possible rational root of 6x^3+23x^2-6x-8 ?

hartnn (hartnn):

getting this ? or went completely overhead? :P

OpenStudy (anonymous):

Im soo lost omg. LOL Could you write it out? and maybe I'll understand step by step

OpenStudy (anonymous):

.. 6x^3 + 23x^2 - 6x - 8 = 6x^3 + ( 24 - 1 ) x^2 - ( 2 + 4 ) x - 8 = 6x^3 - x^2 + 24x^2 - 2x - 4x - 8 = (6x^3 - x^2 - 2x) + (24x^2 - 4x - 8) = x (6x^2 - x - 2) + 4 (6x^2 - x - 2) = (x + 4) (6x^2 - x - 2) = (x + 4) (2x + 1) (3x - 2)

OpenStudy (anonymous):

for the second line. Where does the 24 - 1 and 2+4 come from?

hartnn (hartnn):

ok, for example, the equation 2x^2+3x+7 will have a possible rational root of 7/2

OpenStudy (anonymous):

where does that comee from? :(

OpenStudy (anonymous):

7/2

hartnn (hartnn):